OpenStudy (anonymous):

An object, which is at the origin at time , has initial velocity and constant acceleration? No clue how to do this. help please. An object, which is at the origin at time t=0, has initial velocity V0= (-14.0i - 7.0j)m/s and constant acceleration a=(6.0i + 3.0j)m/s^2.. Find the position x where the object comes to rest (momentarily). Express your answer in terms of the unit vectors i and j. r= _____m???? the i's and j's are considered hats so ihat and jhat...

OpenStudy (amistre64):

i assume we can take the acceleration and integrate it into velocity: a(t) = <6,3> v(t) = <(6t+C),(3t+C)>; where v(0) = <-14,-7> v(t) = <(6t-14),(3t-7)> Now when the velocity = 0, the things stops moving. And that happens when t=7/3, now that we have a time we can calculate this for a position as well by integrating v(t) p(t) = < (3t^2-14t+C) , ((3/2)t^2-7t+C) > from here I would have to assume that p(0) = 0 since it is not stated to be anything that we can calculate. As such, this reduced our vector function to: p(t) = < (3t^2-14t) , ((3/2)t^2-7t) > where t = 7/3 p(7/3) = < -49/3 , -49/6 > but i am not quite sure about the end results; hmm ....